Charles throws a basketball vertically upward at a speed of 5 m/s. The ball reached the maximum height then drops freely. If Charles fail to catch the ball on its way back down, find the height of the bounce of the ball (assume the coefficient of restitution between the ball and ground be 0.6).
1
Expert's answer
2021-10-11T16:20:16-0400

Use the equation of the coefficient of restitution;

coefficient of restitution =velocity of separation"\/" velocity of approach

"0.6=v\/u=v\/5"

"v=3m\/s"

Apply energy conservation after collision;

"1\/2mv^2=mgh"

"1\/2 v^2=gh"

"h=v^2/2g=3^2/(2\times 9.8)=0.46m"